# Sphere Shaping with Dispersion-Aware Sequence Selection for Long-Haul Transmission Systems

### 21 Dec 2023 v1

@@ III E-ESS and D-E-ESS design
$\Psi\overset{\Delta}{=}\frac{\overline{Var}\left[G^{w}\right]}{\overline{E}\left[G^{w}\right]}$ (2)
$G_{i}^{w}\overset{\Delta}{=}\sum_{j=i-w/2}^{i+w/2}\left|X_{j}\right|^{2}$ (3)
$\overline{Var}\left[G^{w}\right]\overset{\Delta}{=}\frac{1}{L_{s}-w}\sum_{i=\frac{w}{2}+1}^{L_{s}-\frac{w}{2}}Var\left[G_{i}^{w}\right]$ (4)
$\overline{E}\left[G^{w}\right]\overset{\Delta}{=}\frac{1}{L_{s}-w}\sum_{i=\frac{w}{2}+1}^{L_{s}-\frac{w}{2}}E\left[G_{i}^{w}\right]$ (5)
Consequently, the assigned signs to the generated amplitudes are not necessary for the evaluation of the metric. After 4D-mapping [28], the sequence selector (EDI calculator) is presented with $2^{\nu n}$ candidate sequences, each with a length of $(l\times n)/4$ 4D-QAM symbols. The sequence selector then identifies and transmits the sequence with the lowest 4D-energy EDI.
To further boost the performance of sequence selection over long-haul links, we introduce sequence selection ESS based on D-EDI. Within the sequence selector, the EDI of dispersed sequences, abbreviated as D-EDI and denoted as $\Psi_{D}$ in Eq. (6), is calculated by averaging the EDI values obtained at multiple points along a linear dispersive fiber without attenuation, as illustrated in Fig. 2. Here, $\mathbf{X}_{m}$ designates the symbol sequence that has propagated over $m\times L_{D}$ kilometers of a linear dispersive fiber without attenuation. $\mathfrak{D}_{D,x}$ in Eq. (7) represents the chromatic dispersion operator of a fiber with a dispersion coefficient $D$ in ps/nm/km and length $x$ km.
@@
$\Psi_{D}\overset{\Delta}{=}\frac{1}{m_{D}+1}\sum_{m=0}^{m_{D}}\Psi\left[\mathbf{X}_{m}\right].$ (6)
$\mathbf{X}_{m}\overset{\Delta}{=}\mathfrak{D}_{D,m\times L_{D}}\left[\mathbf{X}\right].$ (7)
Dispersion is digitally applied over the symbol sequence sampled at one sample per symbol. D-EDI is a metric that requires knowledge of sign bits to compute the form of a dispersed sequence. EDI is the special case of D-EDI where $m_{D}=0$. The sign bits can be fixed through a multi-block FEC-independent sequence selection process as illustrated in Sec. II-B. In this scenario, the sign bits are chosen in two parts:a part from the information stream of the current D-E-ESS block, denoted s, and a part from the FEC output parity bits of the previous D-E-ESS block, denoted c. Once the QAM sequence is selected, it is demapped into a bit sequence and sent into a FEC encoder after discarding the parity bits from the previous D-E-ESS block, c, to generate parity bits for the next D-E-ESS block.
@@
For a single-span link, one may apply D-EDI by cutting the span into multiple sections and adding up EDIs computed over the first sections given that nonlinear distortions are more important at the beginning of each span, where the signal power is still high. In what follows, to keep the computational complexity low for a single-span link, we add up two EDIs, the one computed over the sequences at the transmitter side and the one computed after propagation over the effective length of the fiber span $L_{D\mathrm{(single~{}span)}}=L_{eff}=\frac{1-\mathrm{exp}(-\alpha L)}{\alpha}$ [29] where $\alpha$ is the fiber attenuation in $\textrm{km}^{-1}$ and $L$ is the length of the fiber span in km. For a multi-span link of $m$ spans of length $L$ km each, we may add up the EDIs computed over sequences at the beginning of each span. For the sequence at the beginning of the last span, we apply dispersion corresponding to $m_{D}L_{D}=(m-1)L$ km.
@@
Both E-ESS and D-E-ESS increase the rate loss in each dimension by $\frac{\nu}{l}$ as in [23], which is an extra rate loss term that accounts for the effect of the flipping bits. The total rate loss $R_{loss}$ in each dimension is:
$R_{loss}=H(P_{a})-\frac{k-\nu}{l}.$ (8)
@@
where $H(P_{a})$ is the entropy linked to the distribution $P_{a}$ and $P_{a}$ is the marginal distribution of the amplitudes of the PAM symbols. As the value of $\nu$ increases, the number of candidate sequences also increases, resulting in higher nonlinear shaping gain. However, this comes at the cost of increased rate loss. As a result, an efficient sequence selection mechanism should strike a balance between linear and nonlinear shaping gains. On another note, the utilization of prefix flipping bits eliminates the need for look-up tables (LUTs) and XOR operations proposed in the bit scrambling method in [27], hence simplifying the implementation complexity and avoiding the additional memory requirements of LUTs. Finally, D-EDI computation has clearly a much lower complexity compared to the SSFM-based average NLI metric in [27].
@@
Figure 4: Simulation in single channel, $8$ DMB subcarriers, $110$ GBaud transmission over $30\times 80$ km SSMF link, with a net rate of $880$ Gbit/s. Sequences A and B are two randomly generated sequences over the $4^{\mathrm{th}}$ subcarrier, each consisting of $108$ 4D 64-QAM symbols. Solid lines represent the evolution of EDI at the beginning of each span, dashed lines represent D-EDI computed as an averaged EDI over all spans. D-EDI gives a more accurate estimation of NLI generated by each sequence after transmission.
@@
IV Performance assessment: Single-span transmission
We study the performance of E-ESS and D-E-ESS in two scenarios: first, a single-span transmission scenario similar to the one in [17, 19], and second, a multi-span long-haul transmission scenario using digital-multi-band (DMB) format. Through the two studies, we observe the effects of non-linear distortions over various transmission distances. It is important to remind that these distortions are dominant at higher power levels. Therefore, in a single-span transmission, they primarily occur over the first kilometers of the span. In contrast, in a multi-span scenario with discrete optical amplification, NLI distortions are added at the start of each span.
@@ IV Performance assessment: Single-span transmission
Figure 5: Simulation of $50$ GBaud E-ESS transmission over $205$ km single span SSMF link, $5$ WDM channels, $55$ GHz channel spacing, net rate per channel $400$ Gbit/s. (a): Electrical SNR versus the EDI window length $w$ used in the sequence selector to evaluate the EDI of sequences. $\nu=3$ flipping bits and a single 1D-DM ($n=1$) is used. The power per channel is $9$ dBm. (b): Electrical SNR and GMI at $9$ dBm versus $\log_{2}(N_{S_{108}})$ when cascading 1, 2 or 4 DMs, where $N_{S_{108}}$ is the number of required sequence selections to generate $108$ 4D-symbols. $E\text{-}ESS_{1}^{3}$ exhibits the highest performance in terms of GMI.
We conduct a performance evaluation of the E-ESS and D-E-ESS systems using PDM 64-QAM modulation with a block length of $l=108$ to align with the conditions described in [17, 19]. Our transmitted signal comprised $5$ Wavelength-Division Multiplexing (WDM) channels, each operating at $50$ GBaud, resulting in a raw data rate of $600$ Gbit/s per channel. For all the tested ESS schemes, we employed a Low-Density Parity-Check (LDPC) code with a length of $64800$ bits, following the DVB-S2 standard, and a code rate of $r_{c}=5/6$. Consequently, the achieved net bit rate amounted to $8$ bits per 4D symbol, resulting in a net bit rate of $400$ Gbit/s per channel. The channel spacing is set to $55$ GHz, and we apply root-raised cosine (RRC) pulse shaping with a roll-off factor of $0.1$ to each channel. Standard single-mode fiber (SSMF) is simulated with an attenuation coefficient $\alpha=0.2~{}\mathrm{dB/km}$, chromatic dispersion coefficient $D=17~{}\mathrm{ps/nm/km}$, polarization mode dispersion (PMD) of $0.04~{}\mathrm{ps/\sqrt{km}}$, and a nonlinear parameter $\gamma=1.3~{}\mathrm{(W\cdot km)^{-1}}$ at $\lambda=1550~{}\mathrm{nm}$. Our simulated channel consists in a single-span transmission over a distance of $205~{}\mathrm{km}$, to perform a consistent comparison with [17, 19] in which the Kurtosis-limited (K-ESS) and band-limited (B-ESS) sphere shaping were introduced. We do not add any laser phase noise in the simulation. Following propagation, the central channel underwent several signal processing steps, including optical filtering, matched filtering, chromatic dispersion compensation, and genie-aided multiple-input-multiple-output (MIMO) channel equalization. Subsequently, we applied a fully-data-aided phase filter with a window averaging over $64$ symbols to compensate for phase rotation induced by cross-phase modulation (XPM) as carrier phase recovery (CPR), as elaborated in [30]. Finally, we measured the electric SNR denoted $\mathrm{SNR_{elec}}$ from the equalized constellations and computed the generalized mutual information (GMI) using Eq. (10) from [31].
@@
Figure 6: Simulation of conventional ESS, K-ESS, B-ESS, $E\text{-}ESS_{1}^{3}$, and $D\text{-}E\text{-}ESS_{1}^{4}$ (best-performing scheme) with D-EDI computed after applying a single $21$ km dispersion filter. Same system configuration as for Fig. 5. Left: Electrical SNR versus power per channel. Right: GMI versus power per channel. Both $E\text{-}ESS_{1}^{3}$ and $D\text{-}E\text{-}ESS_{1}^{4}$ demonstrate superior performance when compared to state-of-the art schemes.
@@
First, we focus on the E-ESS scheme. In Fig. 5 (a), we plot the $\mathrm{SNR_{elec}}$ for various EDI window lengths $w$ used by the sequence selector. For this test, we used a single 1D-DM ($n=1$) and $\nu=3$ flipping bits and denote the scheme $E\text{-}ESS_{1}^{3}$ (or in general $E\text{-}ESS_{n}^{\nu}$). This setup allows us to select the most suitable sequence from eight possible candidates ($82^{3}=$ 8). We calculate $\mathrm{SNR_{elec}}$ using equalized symbols after transmitting over a single span with a channel power of 9 dBm, which is near the optimal power level. Our findings show that as $w$ increases, the $\mathrm{SNR_{elec}}$ gradually declines. We find that a window length $w=2$ yields the best performance, and is much lower than the channel memory ($762M=$ 76 when considering the total length of the span). Therefore, we proceed with $w=2$ in subsequent single-span simulations. In Fig. 5 (b), we analyze the performance of E-ESS by varying the number of sequentially cascaded DMs $n$, where each DM has $\nu$ flipping bits. As a complexity metric, we designate the number of tested candidate sequences per108108108108 4D-symbols. We choose $4$ DMs as $4$ is the least common multiple between the three studied values of $n=\left\{1,2,4\right\}$. Moreover, having a block length $l=108$ per DM, 4 DMs will generate $108$ 4D symbols. We denote the number of tested sequences as $N_{S_{108}}$ which is given by:
@@
A single DM with a block length $l=108$ produces $27$ 4D-symbols after 4D-mapping. Consequently, when cascading $n=1,2,$ or $4$ DMs, the candidate sequence lengths will be $27,54,$ and $108$ respectively. To achieve a total sequence of $108$ 4D symbols, the E-ESS transmitter needs to be called $\frac{4}{n}$ times, resulting in $4,2,$ and $1$ calls for $n=1,2,$ and $4$ cascaded DMs respectively. We see from Fig. 5 (b) that, as $N_{S_{108}}$ increases, $\mathrm{SNR_{elec}}$ also increases for $n=\left\{1,2,4\right\}$. However, as the number of flipping bits increases, the rate loss likewise increases. The maximum GMI is hence attained when achieving a trade-off between rate loss and nonlinear gain. Specifically, the largest GMI is achieved with a single 1D-DM at $\log_{2}(N_{S_{108}})=5$ ($32$ candidate sequences), corresponding to the $E\text{-}ESS_{1}^{3}$ configuration.
@@
Next, we also evaluated the performance of the D-E-ESS scheme over the same link by computing D-EDI as the sum of the EDI of the sequence at the transmitter output and its EDI after propagating over a linear dispersive fiber without attenuation of length $L_{eff}$ over which nonlinear effects remain significant. $n=1$ and $\nu=4$ ($D\text{-}E\text{-}ESS_{1}^{4}$) yield the best GMI gains through a similar optimization as the one shown in Fig. 5(b) for E-ESS. In Fig. 6 (a), we show $\mathrm{SNR_{elec}}$ versus the launch power per channel for five different ESS schemes. The D-E-ESS, E-ESS, band-limited ESS (B-ESS), and kurtosis-limited ESS (K-ESS) exhibit SNR gains of approximately $0.8$ dB, $0.6$ dB, $1$ dB, and $0.3$ dB respectively, compared to the conventional ESS at a power level of $10$ dBm. Additionally, they demonstrate up to $1$ dB improvement in the optimal launch power. In Fig. 6 (b), we present the results in terms of GMI per 4D symbol. Notably, the D-E-ESS scheme yields the highest GMI. Specifically, at a power level of $10$ dBm, it achieves a GMI increase of $0.15$ bit/4D compared to B-ESS and a $0.3$ bit/4D improvement compared to conventional ESS. The E-ESS scheme also delivers a GMI gain of $0.1$ bit/4D (respectively $0.2$ bit/4D) in comparison to B-ESS (respectively conventional ESS). It is important to mention that both EDI and D-EDI based shaping approaches exhibit nearly the same linear performance as unconstrained ESS due to their small rate loss. E-ESS and D-E-ESS exhibit similar performance largely because the initial shape of the waveform at the start of the link plays a crucial role in the generation of NLI for single span links. However, a slight performance improvement is achievable by selecting sequences that also exhibit low EDI after the effective length, $L_{eff}$. In terms of rate loss, D-E-ESS demonstrates superiority over B-ESS and K-ESS, with a rate loss of $0.26$ bit/4D compared to $0.5$ bit/4D and $0.27$ bit/4D respectively. The rate losses are not discernible in Fig. 6 (b) between $5$ dBm and the optimal power level because non-linear gains are already occurring within these power ranges. D-E-ESS also outperforms E-ESS in the nonlinear regime. Overall, it outperforms B-ESS in both linear and nonlinear regimes. Finally, we remind that D-E-ESS requires a slightly higher complexity compared to E-ESS since it necessitates knowledge of sign bits and computation of dispersed sequences at the transmitter side.
V Performance assessment: Multi-span transmission
We now evaluate the potential advantages of our schemes when applied to higher baud rate long-haul links employing digital-multi-band (DMB) signals known for their superior resistance to Kerr effects compared to single-carrier schemes [32, 33]. The baud rate in this scenario is set at $110$ Gbaud, utilizing $8$ subcarriers in a single-wavelength configuration. The transmission channel is $30\times 80$ km standard single-mode fiber (SSMF) spans with EDFA amplification. The achieved net data rate per channel is $880$ Gbit/s. The digital signal processing (DSP) blocks employed in this scenario are consistent with those detailed in the preceding section.
V-A Optimization and Transmission Performance
In Fig. 7, we evaluate the performance of our D-E-ESS scheme for different cascaded DMs and different number of flipping bits. In long-distance links, accumulated chromatic dispersion becomes apparent as its induced inter-symbol interference involves hundreds of symbols. For these links, it becomes more important to cascade multiple 1D-DMs to generate extended candidate sequences over which we will compute D-EDI. By doing so, we account for the impact of the accumulated channel memory on the waveform, thus making a better selection process. In addition, for the same complexity level as a single DM, this approach helps in reducing rate loss. As we did in the previous section, the complexity of the scheme is evaluated as $N_{S_{108}}$, the number of tested candidate sequences for the generation of $108$ 4D-symbols. We calculate D-EDI by summing up EDI values of the sequences at the input of each of the $30$ spans constituting the $30\times 80$ km link ($w=2$ corresponding to the memory of a $13.75$ Gbaud transmission over a single $80$ km span). We observe that, as $N_{S_{108}}$ increases, $\mathrm{SNR_{elec}}$ also increases. However, as the number of flipping bits per DM increases, the rate loss increases too. In the scenario of long-haul transmissions, the SNR gain is less significant compared to single-span transmissions. Consequently, the rate loss has a greater impact on the final performance. The trade-off between rate loss and nonlinear gain shows that the largest GMI occurs in the configuration where $n=4$ 1D-DMs are arranged in series, and each 4D-DM has $\nu=1$ flipping bit, denoted $D\text{-}E\text{-}ESS_{4}^{1}$, and corresponding to $\log_{2}(N_{S_{108}})=4$ (i.e. $16$ candidate sequences). The GMI curve for 1DM exhibits a significantly steeper slope compared to the one for 4DM, as the rate loss increases at a slower pace in the case of cascading 4 DMs.

### 05 Mar 2024 v2

@@ III From EDI to D-EDI
To further capture the energy variations within a finite-length sequence propagating over a link, we introduce the EDI of dispersed sequences, abbreviated as D-EDI and defined as the average of EDI values computed at multiple points along a linear dispersive and lossless fiber. This is shown in Fig. 1 and in Eq. (5) in which D-EDI is denoted as $\Psi_{D}$ and $\mathbf{X}^{(N)}$ designates the symbol sequence that has propagated over $N\times L_{D}$ kilometers of an ideal dispersive fiber obtained using Eq. (6) and (7), $\mathrm{m_{D}}$ is the maximum number of dispersion operations and $\Psi\left[\cdot\right]$ is the EDI operator defined in Eq. (1).
$\Psi_{D}\overset{\Delta}{=}\frac{1}{m_{D}+1}\sum_{N=0}^{m_{D}}\Psi\left[\mathbf{X}^{(N)}\right]$ (5)
$\mathbf{X}^{(N)}=\mathfrak{D}_{D,N\times L_{D}}\left[\mathbf{X}\right]$ (6)
@@
$\mathfrak{D}_{D,z}\left[\mathbf{X}\right]\overset{\Delta}{=}\begin{bmatrix}\mathfrak{F}^{-1}\left(\mathfrak{F}\left(\mathbf{x}_{\mathrm{pol_{1}}}\right)\odot\exp\left(j\frac{D\lambda^{2}\pi\boldsymbol{\mathbf{\Delta f}}^{2}z}{c}\right)\right)\\ \mathfrak{F}^{-1}\left(\mathfrak{F}\left(\mathbf{x}_{\mathrm{pol_{2}}}\right)\odot\exp\left(j\frac{D\lambda^{2}\pi\boldsymbol{\mathbf{\Delta f}}^{2}z}{c}\right)\right)\end{bmatrix}$ (7)
The dispersion operator in Eq. (7) consists in a chromatic dispersion filter applied in the frequency domain and capturing the dispersion of a fiber with dispersion coefficient $D$ and length $z$ km. $\mathfrak{F}$ and $\mathfrak{F}^{-1}$ are the fast Fourier transform (FFT) and inverse fast Fourier transform respectively, $\odot$ the element-wise multiplication, $j=\sqrt{-1}$ the imaginary unit, $\lambda$ the central wavelength of the signal, $c$ the speed of light in vacuum and $\mathbf{\Delta f}$ the vector of frequency bins over which the FFT is computed. Dispersion is digitally applied over the perfectly-synchronized unfiltered symbol sequence sampled at one sample per symbol. We clearly see that, unlike EDI, D-EDI requires the knowledge of sign bits to compute the form of a dispersed sequence. EDI is its special case when $m_{D}=0$.
@@
In a single-span link, the signal power is substantially high at the start of the span, leading to significant NLI. To capture waveform variations within the first few kilometers over which the power loss is relatively minimal, we can compute the EDI multiple times shortly after the transmission begins within this span to derive the D-EDI. In what follows, to keep the computational complexity low, for a single-span link, we add up two EDIs, the one computed over the sequences at the transmitter output and the one computed after propagation over the effective length of the fiber span $L_{D\mathrm{(single~{}span)}}=L_{\mathrm{{eff}}}=\frac{1-\mathrm{exp}(-\alpha L)}{\alpha}$ [33] where $\alpha$ is the fiber attenuation in $\textrm{km}^{-1}$ and $L$ is the length of the fiber span in km.
For a multi-span link consisting of $m$ spans of $L$ kilometers each, we may compute EDI several times within each span to capture the evolution of the waveform distortion. In this paper, to limit the complexity of the metric computation, we calculate the EDI of the sequence at the beginning of each span and combine them to obtain the D-EDI, hence $m_{D}=m-1$ and $L_{D}=L$. We may reduce the complexity by computing EDIs less often, i.e. reducing $m_{D}$ and increasing $L_{D}$ as shown in Sec. VI-B.
IV Implementation of EDI and D-EDI: E-SS and D-SS design
Figure 1: Block diagram of FEC independent D-SS in the PAS transmitter: EDI of dispersed sequences (D-EDI) is determined by averaging the calculated EDI at multiple locations along an ideal dispersive fiber. The dispersion is applied at $1$ sample per symbol. EDI is a special case of D-EDI where $m_{D}=0$. The sign bits are fixed through a multi-block FEC-independent sequence selection process as shown in [27].
Figure 2: Structure of cascaded distribution matcher (DM). The output amplitude sequences of $n$ DMs are concatenated to form a longer sequence.
The fundamental concept behind E-SS and D-SS draws inspiration from the list-CCDM [22]. Our approach extends the concept for the use of various distribution matchers (DMs) and we concatenate the output of several DMs to generate a longer candidate sequence. In this study, we only use ESS for distribution matching.
In Fig. 1, we plot the encoding process where input bits are fed into a 1D ESS distribution matcher (DM) with block length $l$ that generates $l$ amplitudes. The input bit stream is composed of two parts: $k-\nu$ information bits and $\nu$ prefix flipping bits. Altering these $\nu$ prefix flipping bits can cause pronounced changes in the indices of the sequences during the ESS encoding process. Such changes are reflected as considerable alterations in the amplitudes across the block. Conveniently, these flipping bits are discarded after the decoding process at the receiver side. The outputs from $n$ 1D-DM can be sequentially concatenated, from head-to-tail, to create extended sequences, which can then undergo a selection process as shown in Fig. 2. This resulting sequence will have a length of $ln$. Cascading DMs have two advantages: they permit obtaining more candidate sequences while maintaining the same rate loss, and at the same time, we maintain the use of short-block-length DMs that showed a clear advantage against long-block-length DM with respect to nonlinear effects. For an E-SS (also true for D-SS) with $n$ cascaded DMs, where each DM includes $\nu$ flipping bits, we can generate $2^{\nu n}$ different candidate sequences. Subsequently, the $ln$ amplitudes are mapped to $L_{s}=ln/4$ 4D-QAM symbols after 4D-mapping [28].
E-SS employs EDI as a metric for sequence selection, hence the assigned signs to the generated amplitudes are not necessary for the evaluation of EDI. The sequence selector then identifies the $2^{\nu n}$ different candidate sequences and transmits the sequence with the lowest 4D-energy EDI.
D-SS employs D-EDI as a metric for sequence selection, which requires the assigned signs for the metric evaluation and selection. The sign bits can be fixed through a multi-block FEC-independent sequence selection process as proposed in [27]. In this scenario, the sign bits are chosen in two parts: a part from the information stream of the current D-SS block, denoted s, and a part from the FEC output parity bits of the previous D-SS block, denoted c as shown in Fig. 1. Once the QAM sequence is selected, it is demapped into a bit sequence and sent into an FEC encoder, after discarding the parity bits from the previous D-SS block c, to generate parity bits for the next D-SS block.
@@ IV Implementation of EDI and D-EDI: E-SS and D-SS design
The sequence selection process introduces an additional rate loss [25, 26], which is $\frac{\nu}{l}$ accounting for the effect of the flipping bits as in [23]. The total rate loss $R_{\mathrm{{loss}}}$ of E-SS and D-SS in each dimension is:
$R_{\mathrm{loss}}=H(P_{a})-\frac{k-\nu}{l}.$ (8)
@@
where $H(P_{a})$ is the entropy linked to the distribution $P_{a}$ and $P_{a}$ is the marginal distribution of amplitudes of 1D PAM symbols. As $\nu$ increases, the number of candidate sequences also increases, resulting in higher nonlinear shaping gain. However, this comes at the cost of an increased rate loss. As a result, an efficient sequence selection mechanism should strike a balance between linear and nonlinear shaping gains.
@@
Figure 3: Simulation in single channel, $8$ DMB subcarriers, $110$ GBaud transmission over $30\times 80$ km SSMF link, with a net rate of $880$ Gbit/s. Sequences A and B are two random sequences generated by single ESS with block length $l=108$ over the $4^{\mathrm{th}}$ subcarrier (i.e. central subcarrier), each consisting of $108$ 4D 64-QAM symbols. Solid lines represent the evolution of EDI at the beginning of each span, dashed lines represent D-EDI computed as an averaged EDI over all spans. D-EDI gives a more accurate estimation of NLI generated by each sequence after transmission.
@@
V Performance assessment: Single-span transmission
@@
Figure 4: Simulation of $50$ GBaud E-SS transmission over $205$ km single span SSMF link, $5$ WDM channels, $55$ GHz channel spacing, net rate per channel $400$ Gbit/s. (a): Electrical SNR versus the EDI window length $w$ used in the sequence selector to evaluate the EDI of sequences. $\nu=3$ flipping bits and a single 1D-DM ($n=1$) is used. The power per channel is $9$ dBm. (b): Electrical SNR and GMI at $9$ dBm versus $\log_{2}(N_{S_{108}})$ when cascading 1, 2 or 4 DMs, where $\mathrm{N_{S_{108}}}$ is the number of required sequence selections to generate $108$ 4D-symbols. $\mathrm{E\text{-}SS_{1}^{3}}$ exhibits the highest performance in terms of GMI (ESS without SS exhibits $13.27$ dB electrical SNR value and $8.46$ bits/4D GMI value).
@@
Figure 5: Simulation of ESS without SS, K-ESS, B-ESS, $\mathrm{E\text{-}SS_{1}^{3}}$, and $\mathrm{D\text{-}SS_{1}^{4}}$ (best-performing scheme) with D-EDI computed after applying a single $21$ km dispersion filter. Same system configuration as for Fig. 4. Left: Electrical SNR versus power per channel. Right: GMI versus power per channel. Both $\mathrm{E\text{-}SS_{1}^{3}}$ and $\mathrm{D\text{-}SS_{1}^{4}}$ demonstrate superior performance when compared to state-of-the art schemes.
@@
The single-span transmission scenario was introduced in Sec. II. First, we focus on the E-SS scheme. In Fig. 4 (a), we plot the $\mathrm{SNR_{elec}}$ for various EDI window lengths $w$ used by the sequence selector. For this test, we used a single 1D-DM ($n=1$) and $\nu=3$ flipping bits and denote the scheme $\mathrm{E\text{-}SS_{1}^{3}}$ (or in general $\mathrm{E\text{-}SS_{n}^{\nu}}$). This setup allows us to select the most suitable sequence from eight possible candidates ($82^{3}=$ 8). We calculate $\mathrm{SNR_{elec}}$ using equalized symbols after transmitting over a single span with a channel power of $9$ dBm, which is near the optimal power level. Our findings show that as $w$ increases, the $\mathrm{SNR_{elec}}$ gradually declines. We find that a window length $w=2$ yields the best performance, and is much lower than the channel memory ($762M=$ 76 when considering the total length of the span) as the NLI is generated mostly at the beginning of the single-span transmission. Therefore, we proceed with $w=2$ in subsequent single-span simulations.
In Fig. 4 (b), we analyze the performance of E-SS by varying the number of sequentially cascaded DMs $n$, where each DM has $\nu$ flipping bits. A single DM with a block length $l=108$ produces $27$ 4D-symbols after 4D-mapping. When cascading $n=1,2$ or $4$ DMs, the candidate sequence lengths will be $27,54$ and $108$ respectively. As a complexity metric, we designate the number of tested candidate sequences per $4$ DMs, i.e. per $108$ 4D-symbols as $\mathrm{N_{S_{108}}}$. We choose $4$ DMs as $4$ is the least common multiple between the three studied values of $n=\left\{1,2,4\right\}$. To achieve a total sequence of $108$ 4D symbols, the E-SS transmitter needs to be called $\frac{4}{n}$ times, resulting in $4,2$ and $1$ calls for $n=1,2,$ and $4$ cascaded DMs respectively. Hence, $\mathrm{N_{S_{108}}}$ is given by:
@@
We see from Fig. 4 (b) that, as $\mathrm{N_{S_{108}}}$ increases, $\mathrm{SNR_{elec}}$ also increases for $n=\left\{1,2,4\right\}$. However, as the number of flipping bits increases, the rate loss likewise increases. The maximum GMI is hence attained when achieving a trade-off between rate loss and nonlinear gain. Specifically, the largest GMI is achieved with a single 1D-DM at $\log_{2}(N_{S_{108}})=5$ ($32$ candidate sequences), corresponding to the $\mathrm{E\text{-}SS_{1}^{3}}$ configuration.
@@
Next, we also evaluated the performance of the D-SS scheme over the same link by computing D-EDI as the sum of the EDI of the sequence at the transmitter output and its EDI after propagating over a linear dispersive fiber without attenuation of length $L_{\mathrm{{eff}}}$ over which nonlinear effects remain significant. $n=1$ and $\nu=4$ ($\mathrm{D\text{-}SS_{1}^{4}}$) yields the best GMI gains through a similar optimization as the one shown in Fig. 4(b) for E-SS. In Fig. 5 (a), we show $\mathrm{SNR_{elec}}$ versus the launch power per channel for five different ESS schemes. The D-SS, E-SS, band-limited ESS (B-ESS), and kurtosis-limited ESS (K-ESS) exhibit SNR gains of approximately $0.8$ dB, $0.6$ dB, $1$ dB, and $0.3$ dB respectively, compared to the ESS without sequence selection (SS) at a power level of $10$ dBm. Additionally, they demonstrate up to $1$ dB improvement in the optimal launch power. In Fig. 5 (b), we present the results in terms of GMI per 4D symbol. Notably, the D-SS scheme yields the highest GMI. Specifically, at a power level of $10$ dBm, it achieves a GMI increase of $0.15$ bits/4D compared to B-ESS and a $0.3$ bits/4D improvement compared to ESS without SS. The E-SS scheme also delivers a GMI gain of $0.1$ bits/4D (respectively $0.2$ bits/4D) in comparison to B-ESS (respectively ESS without SS). It is important to mention that both EDI and D-EDI based shaping approaches exhibit nearly the same linear performance as unconstrained ESS due to their small rate loss. E-SS and D-SS exhibit similar performance largely because the initial shape of the waveform at the start of the link plays a crucial role in the generation of NLI for single span links. However, a slight performance improvement is achievable by selecting sequences that also exhibit low EDI after the effective length $L_{\mathrm{{eff}}}$. In terms of rate loss, D-SS demonstrates superiority over B-ESS and K-ESS, with a rate loss of $0.26$ bits/4D compared to $0.5$ bits/4D and $0.27$ bits/4D respectively. The rate losses are not discernible in Fig. 5 (b) between $5$ dBm and the optimal power level because nonlinear gains are already occurring within these power ranges. D-SS also outperforms E-SS in the nonlinear regime. Overall, it outperforms B-ESS in both linear and nonlinear regimes. Finally, we remind that D-SS requires a slightly higher complexity compared to E-SS since it necessitates knowledge of sign bits and computation of dispersed sequences at the transmitter side.
VI Performance assessment: Multi-span transmission
VI-A Optimization and Transmission Performance